# Qingdao, Shandong 266237, China

### v1

@@ 4.2 The quark-quark correlators
Since the correlation function is a $4\times 4$ matrix in the Dirac space, it can be decomposed by a basis of 16 Dirac matrices. By taking $P^{+}$ as the dominant momentum component of the hadron, the leading-twist (twist-2) PDFs can be projected from the correlation function using the Dirac matrices $\gamma^{+}$, $\gamma^{+}\gamma_{5}$, and $i\sigma^{i+}\gamma_{5}$, while the twist-3 PDFs can be projected by $\gamma^{i}$, $\bm{1}$, $i\gamma_{5}$, $i\sigma^{ij}\gamma_{5}$, and $i\sigma^{-+}\gamma_{5}$. For a spin-1 hadron, the quark-quark correlation function in terms of the TMD PDFs contains the unpolarized, vector-polarized, and tensor-polarized terms. Among them, the unpolarized and vector-polarized parts are identical to those defined for the spin-1/2 hadron. The tensor-polarized part of the quark-quark correlation function, denoted as $\Phi(x,p_{T};T)$, was first parametrized in terms of ten twist-2 TMD PDFs in ref. Bacchetta:2000jk , and subsequently decomposed into 20 twist-3 TMD PDFs, and ten twist-4 TMD PDFs, as shown in ref. Kumano:2020ijt . Up to the subleading twist, a complete decomposition of the tensor-polarized dependent correlation function for a spin-1 hadron is given by
@@
$\Phi^{[i\sigma^{ij}\gamma_{5}]}(x,p_{T};T)=$ $\frac{M}{P^{+}}\bigg{[}-h_{LL}\epsilon_{T}^{ij}S_{LL}-h_{LT}\epsilon_{T}^{ij}\frac{S_{LT}\cdot p_{T}}{M}+h_{TT}\epsilon_{T}^{ij}\frac{p_{T}\cdot S_{TT}\cdot p_{T}}{M^{2}}\bigg{]}$ (67)
@@
Integrating the TMD PDFs in eq. (58) over the transverse momentum, one can obtain three collinear PDFs for tensor-polarized states up to twist-3,
@@
The decomposition of the fragmentation correlation function $\Delta$ for tensor-polarized states can be directly obtained from $\Phi$ in eq. (58) by the kinematic variables replacement
@@
4.3 The quark-gluon-quark correlators
@@
where $\alpha$ is a transverse indice and so forth in the following. The functions with a tilde are interaction-dependent twist-3 distribution functions, which depend on $x$ and $p_{T}^{2}$. The terms $\tilde{f}_{LT}^{\prime}$, $\tilde{g}_{LT}^{\prime}$, $\tilde{f}_{TT}^{\prime}$, and $\tilde{g}_{TT}^{\prime}$ have the same definition as defined in eq. (73). Note that the decomposition of the quark-gluon-quark correlator for tensor-polarized states is similar to that for a vector-polarized hadron, with the distinction arising from the parity-invariant spin tensor. The explicit parametrization of the last term in the curly brackets is not given, as it is irrelevant to the calculation of the hadronic tensor in the SIDIS process.
@@
where $m$ is the mass of the quark. Using the identities $\mathcal{P}_{+}\Phi_{4}=\mathcal{P}_{-}\Phi_{2}=0$ and $\mathcal{P}_{+}\tilde{\Phi}_{A,5}^{\alpha}=\mathcal{P}_{-}\tilde{\Phi}_{A,3}^{\alpha}=0$, we can project out the good component of eq. (98) and express the corresponding correlators as
@@
Substituting the expression of quark-quark correlator (58) containing $\Phi_{2}$ and $\Phi_{3}$ and quark-gluon-quark correlator $\tilde{\Phi}_{A}^{\alpha}$ (92) into eq. (100), one can establish the relations between the twist-2 and twist-3 functions. For a tensor-polarized spin-1 hadron, we derive 12 relations between T-even functions,
@@
where we kept the sign of the functions with a tilde in eq. (92) consistent with those defined in quark-quark correlator (58). The functions with a tilde in eqs. (101)-(124) vanish in the so-called Wandzura–Wilczek approximation, which is equivalent to neglecting the quark-gluon-quark correlator (see, e.g., refs. Wandzura:1977qf ; Accardi:2009au ; Bastami:2018xqd ). We can find that the $S_{LL}$-dependent relations share the same forms as the unpolarized ones in ref. Bacchetta:2006tn . Each of the TMD PDFs defined in eq. (58) have the corresponding relations derived from the equation of motion, except for four additional ones specified by the identity (73).
@@
4.4 The results for structure functions
@@
Up to twist-3, only two structure functions are zero, given by eqs. (134) and (139). However, if we limit ourselves to the twist-2 level, 12 out of the 23 structure functions vanish. We should also note that structure functions involving twist-2 terms exhibit dependence on an odd number of azimuthal angles ($\phi_{h}$, $\phi_{LT}$, $\phi_{TT}$), whereas twist-3 structure functions depend on an even number of azimuthal angles. These nonvanishing structure functions can be utilized to study the tensor-polarized TMD PDFs for a spin-1 particle.
@@
$b_{1}$ $=-\frac{1}{2(\gamma^{2}+1)}\Big{(}F_{U(LL),T}+F_{U(TT)}^{\cos(2\phi_{TT})}\Big{)},$ (164)
$b_{2}$ $=-\frac{x_{d}}{(\gamma^{2}+1)^{2}}\Big{[}F_{U(LL),L}+F_{U(LL),T}+2\gamma F_{U(LT)}^{\cos\phi_{LT}}+(2\gamma^{2}+1)F_{U(TT)}^{\cos(2\phi_{TT})}\Big{]},$ (165)
$\gamma^{2}b_{3}$ $=\frac{x_{d}}{3(\gamma^{2}+1)^{2}}\Big{[}-\gamma^{2}F_{U(LL),L}-\gamma^{2}F_{U(LL),T}-2\gamma^{3}F_{U(LT)}^{\cos\phi_{LT}}$
@@
$\gamma^{2}b_{4}$ $=-\frac{x_{d}}{3(\gamma^{2}+1)^{2}}\Big{[}\gamma^{2}F_{U(LL),L}+\gamma^{2}F_{U(LL),T}-2\gamma(2\gamma^{2}+3)F_{U(LT)}^{\cos\phi_{LT}}$
$\qquad+(2\gamma^{4}+13\gamma^{2}+12)F_{U(TT)}^{\cos(2\phi_{TT}}\Big{]},$ (167)
where the prefactor $\gamma^{2}$ of $b_{3}$ and $b_{4}$ indicate that they are twist-4 structure functions Hoodbhoy:1988am . The measurement of $b_{1}$ can reveal how nuclear interaction affects the collinear partonic structure of spin particles. Recently, an experiment has been proposed at JLab to measure the deuteron tensor structure function $b_{1}$ by deep inelastic scattering with a tensor-polarized deuterium target Poudel:2025tac ; Poudel:2025nof . The tensor structure can also be investigated on a polarized proton-deuteron Drell-Yan process at Fermilab Keller:2022abm ; Keller:2020wan . Therefore, one can in principle extract the collinear PDFs $f_{1LL}(x)$ and $f_{LT}(x)$ from future experiment data.
@@ Acknowledgements.
Appendix A The hadronic tensor in terms of partonic functions

### v2

@@ 4.2 Quark-quark correlators
Since the correlation function is a $4\times 4$ matrix in the Dirac space, it can be decomposed by a basis of 16 Dirac matrices. By taking $P^{+}$ as the dominant momentum component of the hadron, the elements in the correlation functions are ordered in powers of $M/P^{+}$. Each additional power of $M/P^{+}$ will lead to a power of $M/Q$ suppressed contribution to the cross section Levelt:1993ac . Here the twist $t$ of TMD PDFs or FFs is associated to $(M/Q)^{t-2}$ in structure functions Mulders:1995dh . The leading-twist (twist-2) PDFs can be projected from the correlation function using the Dirac matrices $\gamma^{+}$, $\gamma^{+}\gamma_{5}$, and $i\sigma^{i+}\gamma_{5}$, while the twist-3 PDFs can be projected by $\gamma^{i}$, $\bm{1}$, $i\gamma_{5}$, $i\sigma^{ij}\gamma_{5}$, and $i\sigma^{-+}\gamma_{5}$. For a spin-1 hadron, the quark-quark correlation function in terms of the TMD PDFs contains the unpolarized, vector-polarized, and tensor-polarized terms. Among them, the unpolarized and vector-polarized parts are identical to those defined for the spin-1/2 hadron. The tensor-polarized part of the quark-quark correlation function, denoted as $\Phi(x,p_{T};T)$, was first parametrized in terms of ten twist-2 TMD PDFs in ref. Bacchetta:2000jk , and subsequently decomposed into 20 twist-3 TMD PDFs, and ten twist-4 TMD PDFs, as shown in ref. Kumano:2020ijt . Up to the subleading twist, a complete decomposition of the tensor-polarized dependent correlation function for a spin-1 hadron is given by
@@
$\Phi^{[i\sigma^{ij}\gamma_{5}]}(x,p_{T};T)=$ $\frac{M}{P^{+}}\bigg[-h_{LL}\epsilon_{T}^{ij}S_{LL}-h_{LT}\epsilon_{T}^{ij}\frac{S_{LT}\cdot p_{T}}{M}+h_{TT}\epsilon_{T}^{ij}\frac{p_{T}\cdot S_{TT}\cdot p_{T}}{M^{2}}\bigg],$ (67)
@@
Integrating the TMD PDFs in eq. (58) over the transverse momentum, one can obtain three collinear PDFs for tensor-polarized components up to twist-3,
@@
The decomposition of the fragmentation correlation function $\Delta$ for tensor-polarized components can be directly obtained from $\Phi$ in eq. (58) by the kinematic variables replacement
@@
4.3 Quark-gluon-quark correlators
@@
where $\alpha$ is a transverse indice and so forth in the following. The functions with a tilde are interaction-dependent twist-3 distribution functions, which depend on $x$ and $p_{T}^{2}$. The terms $\tilde{f}_{LT}^{\prime}$, $\tilde{g}_{LT}^{\prime}$, $\tilde{f}_{TT}^{\prime}$, and $\tilde{g}_{TT}^{\prime}$ have the same definition as defined in eq. (73). Note that the decomposition of the quark-gluon-quark correlator for tensor-polarized components is similar to that for a vector-polarized hadron, with the distinction arising from the parity-invariant spin tensor. The explicit parametrization of the last term in the curly brackets is not given, as it is irrelevant to the calculation of the hadronic tensor in the SIDIS process.
@@
where $m$ is the mass of the quark. With the projection operators $\mathcal{P}_{+}=\frac{1}{2}\gamma^{-}\gamma^{+}$ and $\mathcal{P}_{-}=\frac{1}{2}\gamma^{+}\gamma^{-}$ Jaffe:1996zw , we have the identities $\mathcal{P}_{+}\Phi_{4}=\mathcal{P}_{-}\Phi_{2}=0$ and $\mathcal{P}_{+}\tilde{\Phi}_{A,5}^{\alpha}=\mathcal{P}_{-}\tilde{\Phi}_{A,3}^{\alpha}=0$. Using these relations, we can project out the good component of eq. (98) and express the corresponding correlators as
@@
Substituting the expression of the quark-quark correlator (58) containing $\Phi_{2}$ and $\Phi_{3}$ and the quark-gluon-quark correlator $\tilde{\Phi}_{A}^{\alpha}$ (92) into eq. (100), one can establish the relations between the twist-2 and twist-3 functions. For a tensor-polarized spin-1 hadron, we derive 12 relations between T-even functions,
@@
where we kept the sign of the functions with a tilde in eq. (92) consistent with those defined in the quark-quark correlator (58). The functions with a tilde in eqs. (101)-(124) vanish in the so-called Wandzura–Wilczek approximation, which is equivalent to neglecting the quark-gluon-quark correlator (see, e.g., refs. Wandzura:1977qf ; Accardi:2009au ; Bastami:2018xqd ). We can find that the $S_{LL}$-dependent relations share the same forms as the unpolarized ones in ref. Bacchetta:2006tn . Each of the TMD PDFs defined in eq. (58) have the corresponding relations derived from the equation of motion, except for four additional ones specified by the identity (73).
@@
4.4 Results for structure functions
@@
Up to twist-3, only two structure functions are zero, given by eqs. (134) and (139). However, if we limit ourselves to the twist-2 level, 12 out of the 23 structure functions vanish. We should also note that structure functions involving twist-2 terms exhibit dependence on an even number of azimuthal angles ($\phi_{h}$, $\phi_{LT}$, $\phi_{TT}$), whereas twist-3 structure functions depend on an odd number of azimuthal angles. For example, the $F_{U(LT)}^{\cos(2\phi_{h}-\phi_{LT})}$ in eq. (141) depends on $2\phi_{h}$ and $\phi_{LT}$, which sum to three (odd) azimuthal angles, and then it receives the leading contribution at twist-3. The $F_{U(LT)}^{\cos(3\phi_{h}-\phi_{LT})}$ in eq. (143) depends on $3\phi_{h}$ and $\phi_{LT}$, which sum to four (even) azimuthal angles, and then it receives the leading contribution at twist-2. These nonvanishing structure functions can be utilized to study the tensor-polarized TMD PDFs for a spin-1 particle.
@@
$b_{1}$ $=-\frac{1}{2(\gamma^{2}+1)}\Big(\frac{3}{2}F_{U(LL),T}+F_{U(TT)}^{\cos(2\phi_{TT})}\Big),$ (164)
$b_{2}$ $=-\frac{x_{d}}{(\gamma^{2}+1)^{2}}\Big[\frac{3}{2}\big(F_{U(LL),L}+F_{U(LL),T}\big)+4\gamma F_{U(LT)}^{\cos\phi_{LT}}+(2\gamma^{2}+1)F_{U(TT)}^{\cos(2\phi_{TT})}\Big],$ (165)
$\gamma^{2}b_{3}$ $=\frac{x_{d}}{3(\gamma^{2}+1)^{2}}\Big[-\frac{3}{2}\gamma^{2}\big(F_{U(LL),L}+F_{U(LL),T}\big)-4\gamma^{3}F_{U(LT)}^{\cos\phi_{LT}}$
@@
$\gamma^{2}b_{4}$ $=-\frac{x_{d}}{3(\gamma^{2}+1)^{2}}\Big[\frac{3}{2}\gamma^{2}\big(F_{U(LL),L}+F_{U(LL),T}\big)-4\gamma(2\gamma^{2}+3)F_{U(LT)}^{\cos\phi_{LT}}$
$\qquad+(2\gamma^{4}+13\gamma^{2}+12)F_{U(TT)}^{\cos(2\phi_{TT})}\Big].$ (167)
Compared to the SIDIS process off a spin-1/2 target, many new structure functions are introduced for tensor polarized spin-1 target. While all these structure functions are in principle measurable in experiments via corresponding spin asymmetries and the separation of various azimuthal modulations, we would like to discuss and emphasize some particular terms considering experimental feasibility and physical insights.
In general, leading-twist observables are experimentally favorable, because higher-twist contributions are formally suppressed in powers of $1/Q$ and are therefore more challenging. Measurements of the structure functions containing $D_{1}$, such as $F_{U(LL),T}$, $F_{U(LT),T}^{\cos(\phi_{h}-\phi_{LT})}$, $F_{L(LT)}^{\sin(\phi_{h}-\phi_{LT})}$, $F_{U(TT),T}^{\cos(2\phi_{h}-2\phi_{LT})}$, and $F_{L(TT)}^{\sin(2\phi_{h}-2\phi_{TT})}$ should provide the clearest access to tensor-polarized TMDs, as they are convoluted with the relatively well-constrained unpolarized fragmentation function $D_{1}$. The kinematic coverage available at JLab and at the future EIC makes these leading-twist channels especially promising for early measurements.
The $S_{LL}$ sector is expected to be the most accessible experimentally. An $S_{LL}$-polarized deuteron target has already been realized in the HERMES measurement HERMES:2005pon , and a new experiment at JLab has been proposed to measure the deuteron tensor structure function with an $S_{LL}$-polarized target Jlabproposal . The tensor-polarized PDFs and TMDs can probe new feature of the deuteron structure beyond a simple bound system of a proton and a neutron. In particular, the large difference between the HERMES data and the standard deuteron model calculations indicates such a possibility. The structure function $b_{1}$ in inclusive scattering remains a primary aim and is a key measurement of the proposed experiment at JLab Jlabproposal . Its counterpart in SIDIS, $F_{U(LL),T}$, is also of particular interest, and dedicated analyses for extracting $f_{1LL}$ have been proposed Poudel:2025nof ; Poudel:2025tac . Compared to inclusive DIS, the transverse-momentum $P_{h\perp}$ dependence in SIDIS can discriminate among competing models for $b_{1}$, and the convolution with TMD fragmentation functions offers sensitivity to the flavor separation. Besides, the azimuthal dependent term $F_{U(LL)}^{\cos 2\phi_{h}}$, which involves the only other leading-twist $S_{LL}$-dependent TMD PDF $h_{1LL}$, also deserves attention.
For the $S_{LT}$ sector, in addition to the leading-twist terms, the twist‑3 modulation term $F_{U(LT)}^{\cos\phi_{LT}}$ is also noteworthy, because it survives upon integrating over $P_{h\perp}$ and even contributes in inclusive DIS. The $\sin\phi_{LT}$ modulation term is naively T‑odd and should vanish after the integration over $P_{h\perp}$, providing a test on the role of intrinsic transverse momentum and gauge‑link–induced phases in generating T‑odd effects Bacchetta:2002xd .
For a realistic demonstration of experimental feasibility, which relies on the expected magnitude of corresponding asymmetries, model calculations are desired. The traditional convolution model is a commonly used approach for describing the deuteron structure Hoodbhoy:1988am ; Frankfurt:1983qs ; Cosyn:2017fbo ; Khan:1991qk . Although its results differ from those of the HERMES data, one may improve it and perform numerical estimates for the tensor-polarized structure functions, such as by extending it to the partonic level. Additionally, the calculation of the tensor polarized TMD PDFs for the deuteron is underway by utilizing light-front wave functions Sargsian:2022rmq ; Kaur:2025css ; Mondal:2025fdl . These model estimates will provide some quantitative guidance for future experimental measurements of the deuteron structures.
@@ Acknowledgements.
Appendix A Hadronic tensor in terms of partonic functions